OpenStudy (cvvxvxcvxcvxc):

56=2.5x+(12-x)5.75 solve for x

alones (alones):

56=2.5x+(12−x)(5.75) Simplify both sides of the equation. 56=2.5x+(12−x)(5.75) 56=−3.25x+69 Thank just flip the equation −3.25x+69=56

OpenStudy (cvvxvxcvxcvxc):

thx love you

alones (alones):

and subtract 69 from both sides −3.25x+69−69=56−69 −3.25x=−13 divide both sides by -3.25. \(\frac{ −3.25x}{ −3.25 }\)=\(\frac{ −13}{ −3.25 }\) Overall i got 4 for my answer /: since x=4

alones (alones):

Lol that was random. Yw ;)